# OverLoCK: An Overview-first-Look-Closely-next ConvNet with Context-Mixing Dynamic Kernels

### v1

@@ Abstract
In the human vision system, top-down attention plays a crucial role in perception, wherein the brain initially performs an overall but rough scene analysis to extract salient cues (i.e., overview first), followed by a finer-grained examination to make more accurate judgments (i.e., look closely next). However, recent efforts in ConvNet designs primarily focused on increasing kernel size to obtain a larger receptive field without considering this crucial biomimetic mechanism to further improve performance. To this end, we propose a novel pure ConvNet vision backbone, termed OverLoCK, which is carefully devised from both the architecture and mixer perspectives. Specifically, we introduce a biomimetic Deep-stage Decomposition Strategy (DDS) that fuses semantically meaningful context representations into middle and deep layers by providing dynamic top-down context guidance at both feature and kernel weight levels. To fully unleash the power of top-down context guidance, we further propose a novel Context-Mixing Dynamic Convolution (ContMix) that effectively models long-range dependencies while preserving inherent local inductive biases even when the input resolution increases. These properties are absent in previous convolutions. With the support from both DDS and ContMix, our OverLoCK exhibits notable performance improvement over existing methods. For instance, OverLoCK-T achieves a Top-1 accuracy of 84.2%, significantly surpassing ConvNeXt-B while only using around one-third of the FLOPs/parameters. On object detection with Cascade Mask R-CNN, our OverLoCK-S surpasses MogaNet-B by a significant 1% in APb. On semantic segmentation with UperNet, our OverLoCK-T remarkably improves UniRepLKNet-T by 1.7% in mIoU. Code is publicly available at https://bit.ly/4bdbmdl.
@@ 1 Introduction
Top-down neural attention [1, 2, 3] is a crucial perception mechanism in the human vision system, which suggests that the brain initially processes a visual scene to quickly form an overall high-level perception, which goes back to fuse with the sensory input, enabling the brain to make more accurate judgments, such as object locations, shapes, and categories. Many previous works have incorporated such top-down attention into vision models, but some of them are unsuitable for building modern vision backbones due to incompatible model designs [4, 5, 6, 7, 8] while the remaining methods primarily focus on recurrent architectures [9, 10, 11, 12, 13], which introduce additional computational overhead due to recurrent operations, resulting in a suboptimal tradeoff between performance and computational complexity.
@@
A key property of the top-down attention mechanism is to utilize feedback signals as explicit guidance to localize meaningful regions in a scene [1]. However, the classic hierarchical architecture employed in most existing vision backbones [14, 15, 16, 17, 18, 19] contrasts with this biological mechanism as it progressively encodes features from lower to higher levels so that the input features of a layer rely solely on features from previous layers. Hence, there is a lack of explicit top-down semantic guidance in the operations at intermediate layers. To investigate this, we visualize class activation maps [20] and effective receptive fields (ERFs) [21] of three representative hierarchical vision models: Swin-T [15], ConvNeXt-T [18], and VMamba-T [19]. As shown in Figure 2, these image classification models struggle to accurately localize objects with the correct category label in feature maps, especially in Stage 3, which is farther from the classifier layer, despite capturing long-range dependencies in varying degrees. Therefore, how to develop a modern ConvNet that leverages the top-down attention mechanism while achieving an excellent performance-complexity tradeoff remains an open problem.
@@
As a top-down context contains information across the entire input image, to fully unleash its power and absorb its information into convolution kernels, the Focus-Net should utilize a powerful dynamic convolution as the token mixer, capable of adaptively modeling long-range dependencies to produce large receptive fields while preserving local inductive biases to capture nuanced local details. Nonetheless, we find that existing convolutions cannot meet these requirements simultaneously. Unlike self-attention mechanisms [17, 22, 23] and State Space Models [24, 25, 19] that can adaptively model long-range dependencies at various input resolutions, large kernel convolutions [18, 26, 27, 28, 29] and dynamic convolutions [30, 31, 32] are still confined to finite regions due to fixed kernel sizes even when input images have an increasingly large resolution, suggesting a weak long-range modeling ability. Although deformable convolutions [33, 34] can alleviate these issues to a certain extent, a deformable kernel shape sacrifices the inherent inductive bias of convolutions, giving rise to a relatively weak local perception ability. Hence, enabling pure convolutions to possess dynamic global modeling capabilities comparable to those of Transformer- and Mamba-based models while preserving strong inductive biases remains a challenge.
@@ 2 Related Work
Evolution of ConvNets. Since the debut of AlexNet [37], ConvNets gradually became the dominant architecture in computer vision. VGGNet [38] introduced the concept of stacking small kernels to build deep networks. ResNet [14] and DenseNet [39] further proposed skip-connections to address the gradient vanishing/exploding issues in deep networks. However, with the rise of Vision Transformers [40, 41, 16, 15], ConvNets’ dominance in vision tasks has been challenged. Hence, recent methods have proposed increasingly larger kernel sizes to mimic the self-attention mechanism [40] and establish long-range dependencies [18, 26, 27, 36, 29, 28]. ConvNeXt [18] pioneered the use of 7$\times$7 kernels to build vision backbones, surpassing the performance of Swin Transformer [15]. RepLKNet [26] further explored the promising performance of very large kernels by using 31$\times$31 kernels. On the other hand, gated mechanisms have been extensively explored in ConvNets [42, 43, 35, 44, 45]. For instance, MogaNet [35] introduced a multi-order gated aggregation module to enhance the capacity for refining multi-scale feature representations. StarNet [44] unveiled the underlying reason for the superior performance of element-wise multiplications in the gated mechanism. More recently, RDNet [46] rethought the design of DenseNet and proposed an efficient densely-connected ConvNet. Unlike previous work, this paper focuses on improving the performance of ConvNets from both the architectural and mixer perspectives.
@@
Biomimetic Vision Models. The human vision system has inspired the design of a number of excellent vision backbone models. For instance, several advanced vision backbones [42, 52, 36] have been inspired by the peripheral perception mechanism [53], achieving notable performance. Likewise, the top-down attention mechanism [1, 2, 3] has promoted developments in computer vision and machine learning, such as enhancing performance on specific tasks [4, 5, 6, 7, 8], exploring new learning algorithms [54], and designing generic architectures with a recurrent style [9, 10, 11, 12]. Recently, AbsViT [13] introduced a feedback-based Vision Transformer backbone that reuses network outputs to recalibrate early features. In contrast to the above works, we propose a novel modern ConvNet-based vision backbone network that can efficiently generate and utilize top-down guidance, achieving significant performance gains across diverse vision tasks.
@@ 4.1 Image Classification
Results. As shown in Table 3, our pure ConvNet model achieves notable performance improvements over other competitors. For instance, OverLoCK-XT surpasses a strong Transformer-based model (BiFormer-T [23]) and a recent large kernel ConvNet (UniRepLKNet-N [29]) by significant 1.3% and 1.1% in Top-1 accuracy, respectively. For Tiny models, our OverLoCK-T also attains the best performance compared with other methods, achieving 84.2% Top-1 accuracy, which improves upon MogaNet-S [35] and PeLK-T [36] by 0.8% and 1.6% in Top-1 accuracy, respectively. When scaling up to larger models, our OverLoCK still maintains a significant advantage. Specifically, OverLoCK-S improves upon BiFormer-B and UniRepLKNet-S by notable 0.5% and 0.9% in Top-1 accuracy, respectively, with comparable computational complexity. Regarding the largest model, OverLoCK-B achieves an impressive 85.1% Top-1 accuracy, outperforming MaxViT-B by 0.2% in Top-1 accuracy with significantly lower computational complexity.
Meanwhile, we evaluate the throughput of different models using a batch size of 128 on a single NVIDIA L40S GPU. Figure 1 demonstrates that our OverLoCK achieves an excellent tradeoff between speed and accuracy. For instance, OverLoCK-S surpasses MogaNet-B with over 100 imgs/s in throughput, while significantly increasing Top-1 accuracy from 84.3% to 84.8%. Similarly, OverLoCK-XT exceeds BiFormer-T by over 600 imgs/s in throughput, while remarkably improving Top-1 accuracy by 1.3%. Overall, to the best of our knowledge, OverLoCK is the first pure ConvNet model to achieve such substantial performance gains over strong baselines on ImageNet-1K.
@@ Appendix A More Ablation Studies
Impact of Kernel Sizes. We compared the performance under various settings of kernel sizes, as outlined in Table A (the definition of the kernel size in our proposed method is given in Section 3.3). The results indicate that the configuration $\left\{[17,15,13],[7],[13,7]\right\}$ yields the optimal performance on both image classification and semantic segmentation tasks. Further enlarging the kernels does not lead to additional improvements.
@@
Table G: A comparison of image classification with 384×384 inputs.
@@ Appendix C Speed Analysis
We provide a comparison of speed-accuracy trade-off in Figure 1. More details are listed in Table I, where an OverLoCK variant often achieves faster speed and higher accuracy simultaneously than a larger variant of another network, demonstrating an excellent trade-off between speed and accuracy. For instance, OverLoCK-XT achieves 1672 imgs/s in throughput, improving upon Swin-T by over 300 imgs/s, while significantly enhancing Top-1 accuracy by 1.4%. Also, OverLoCK-T achieves about 200 imgs/s improvement in throughput compared to ConvNeXt-B while achieving better performance at the cost of only around one-third of the FLOPS. When compared to more advanced models, OverLoCK still exhibits significant advantages. For example, OverLoCK-S surpasses MogaNet-B by over 100 imgs/s in throughput while increasing Top-1 accuracy from 84.3% to 84.8%. Likewise, OverLoCK-XT surpasses BiFormer-T by over 600 imgs/s in throughput while remarkably improving Top-1 accuracy by 1.3%.

### v2

@@ Abstract
Top-down attention plays a crucial role in the human vision system, wherein the brain initially obtains a rough overview of a scene to discover salient cues (i.e., overview first), followed by a more careful finer-grained examination (i.e., look closely next). However, modern ConvNets remain confined to a pyramid structure that successively downsamples the feature map for receptive field expansion, neglecting this crucial biomimetic principle. We present OverLoCK, the first pure ConvNet backbone architecture that explicitly incorporates a top-down attention mechanism. Unlike pyramid backbone networks, our design features a branched architecture with three synergistic sub-networks: 1) a Base-Net that encodes low/mid-level features; 2) a lightweight Overview-Net that generates dynamic top-down attention through coarse global context modeling (i.e., overview first); and 3) a robust Focus-Net that performs finer-grained perception guided by top-down attention (i.e., look closely next). To fully unleash the power of top-down attention, we further propose a novel context-mixing dynamic convolution (ContMix) that effectively models long-range dependencies while preserving inherent local inductive biases even when the input resolution increases, addressing critical limitations in existing convolutions. Our OverLoCK exhibits a notable performance improvement over existing methods. For instance, OverLoCK-T achieves a Top-1 accuracy of 84.2%, significantly surpassing ConvNeXt-B while using only around one-third of the FLOPs/parameters. On object detection, our OverLoCK-S clearly surpasses MogaNet-B by 1% in APb. On semantic segmentation, our OverLoCK-T remarkably improves UniRepLKNet-T by 1.7% in mIoU. Code is publicly available at https://rb.gy/wit4jh.
@@ 1 Introduction
Top-down neural attention [58, 17, 40] is a crucial perception mechanism in the human vision system, which suggests that the brain initially processes a visual scene to quickly form an overall high-level perception, which goes back to fuse with the sensory input, enabling the brain to make more accurate judgments, such as object locations, shapes, and categories. Many previous works have incorporated such top-down attention into vision models, but some of them are unsuitable for building modern vision backbones due to incompatible model designs [28, 71, 1, 53, 6] while the remaining methods primarily focus on recurrent architectures [3, 76, 4, 54, 60], which introduce additional computational overhead due to recurrent operations, resulting in a suboptimal trade-off between performance and computational complexity.
@@
A key property of the top-down attention mechanism is the use of feedback signals as explicit guidance to locate meaningful regions in a scene [58]. However, the classic hierarchical architecture employed in most existing vision backbones [23, 44, 66, 67, 45, 43] contrasts with this biological mechanism, as it progressively encodes features from lower to higher levels so that the input features of a layer rely solely on features from previous layers. Hence, there is a lack of explicit top-down semantic guidance in the operations at intermediate layers. To investigate this, we visualize the class activation maps [59] and the effective receptive fields (ERFs) [49] of three representative hierarchical vision models. Swin-T [44], ConvNeXt-T [45], and VMamba-T [43]. As shown in Figure 2, these image classification models struggle to accurately localize objects with the correct category label in the feature maps, especially in Stage 3, which is farther from the classifier layer, despite capturing long-range dependencies in varying degrees. Therefore, how to develop a modern ConvNet that leverages the top-down attention mechanism while achieving an excellent performance-complexity trade-off remains an open problem.
@@
As a top-down context contains information across the entire input image, to fully unleash its power and absorb its information into convolution kernels, the Focus-Net should utilize a powerful dynamic convolution as the token mixer, capable of adaptively modeling long-range dependencies to produce large receptive fields while preserving local inductive biases to capture nuanced local details. Nonetheless, we find that existing convolutions cannot meet these requirements simultaneously. Unlike self-attention mechanisms [14, 66, 67, 64, 80] and State Space Models [18, 81, 43, 47, 16] that can adaptively model long-range dependencies at various input resolutions, large kernel convolutions [45, 12, 42, 74, 13] and dynamic convolutions [8, 37, 36] are still confined to finite regions due to fixed kernel sizes even when input images have an increasingly large resolution, suggesting a weak long-range modeling ability. Although deformable convolutions [10, 69] can alleviate these issues to a certain extent, a deformable kernel shape sacrifices the inherent inductive bias of convolutions, giving rise to a relatively weak local perception ability. Hence, enabling pure convolutions to possess dynamic global modeling capabilities comparable to those of Transformer- and Mamba-based models while preserving strong inductive biases remains a challenge.
@@ 2 Related Work
Evolution of ConvNets. Since the debut of AlexNet [33], ConvNets gradually became the dominant architecture in computer vision. VGGNet [61] introduced the concept of stacking small kernels to build deep networks. ResNet [23] and DenseNet [30] further proposed skip-connections to address the gradient vanishing/exploding issues in deep networks. However, with the rise of Vision Transformers [14, 63, 66, 44, 80, 48], ConvNets’ dominance in vision tasks has been challenged. Hence, recent methods have proposed increasingly larger kernel sizes to mimic the self-attention mechanism [14] and establish long-range dependencies [45, 12, 42, 7, 13, 74]. ConvNeXt [45] pioneered the use of 7$\times$7 kernels to build vision backbones, surpassing the performance of Swin Transformer [44]. RepLKNet [12] further explored the promising performance of very large kernels by using 31$\times$31 kernels. On the other hand, gated mechanisms have been extensively explored in ConvNets [73, 56, 39, 50, 51]. For instance, MogaNet [39] introduced a multi-order gated aggregation module to enhance the capacity for refining multi-scale feature representations. StarNet [50] unveiled the underlying reason for the superior performance of element-wise multiplications in the gated mechanism. More recently, RDNet [32] rethought the design of DenseNet and proposed an efficient densely-connected ConvNet. Unlike previous work, this paper focuses on improving the performance of ConvNets from both the architectural and mixer perspectives.
@@
Biomimetic Vision Models. The human vision system has inspired the design of many excellent vision backbone networks. For instance, several advanced vision backbones [73, 52, 7] have been inspired by the peripheral perception mechanism [35], achieving notable performance. Likewise, the top-down attention mechanism [58, 17, 40] has promoted developments in computer vision and machine learning, such as enhancing performance on specific tasks [28, 71, 1, 53, 6], exploring new learning algorithms [77], and designing generic architectures with a recurrent style [3, 76, 4, 54]. Recently, AbsViT [60] introduced a feedback-based Vision Transformer backbone that reuses network outputs to recalibrate early features. In contrast to the above works, we propose a novel modern ConvNet-based vision backbone network that can efficiently generate and utilize top-down guidance, achieving significant performance gains across diverse vision tasks.
@@ 3.2 Dynamic Convolution with Context-Mixing
3.3 Network Architecture
4 Experiments
4.1 Image Classification
Results. As shown in Table 3, our pure ConvNet model achieves notable performance improvements over other competitors. For instance, OverLoCK-XT surpasses a strong Transformer-based model (BiFormer-T [80]) and a recent large kernel ConvNet (UniRepLKNet-N [13]) by significant 1.3% and 1.1% in Top-1 accuracy, respectively. For Tiny models, our OverLoCK-T also attains the best performance compared with other methods, achieving 84.2% Top-1 accuracy, which improves upon MogaNet-S [39] and PeLK-T [7] by 0.8% and 1.6% in Top-1 accuracy, respectively. When scaling up to larger models, our OverLoCK still maintains a significant advantage. Specifically, OverLoCK-S improves upon BiFormer-B and UniRepLKNet-S by notable 0.5% and 0.9% in Top-1 accuracy, respectively, with comparable computational complexity. Regarding the largest model, OverLoCK-B achieves an impressive 85.1% Top-1 accuracy, outperforming MaxViT-B by 0.2% in Top-1 accuracy with significantly lower computational complexity. Meanwhile, we evaluate the throughput of different models using a batch size of 128 on a single NVIDIA L40S GPU. Figure 1 demonstrates that our OverLoCK achieves an excellent tradeoff between speed and accuracy. For instance, OverLoCK-S surpasses MogaNet-B with over 100 imgs/s in throughput, while significantly increasing Top-1 accuracy from 84.3% to 84.8%. Similarly, OverLoCK-XT exceeds BiFormer-T by over 600 imgs/s in throughput, while remarkably improving Top-1 accuracy by 1.3%. Overall, to the best of our knowledge, OverLoCK is the first pure ConvNet model to achieve such substantial performance gains over strong baselines on ImageNet-1K.
@@ Appendix A More Ablation Studies
Impact of Kernel Sizes. We compared the performance under various settings of kernel sizes, as outlined in Table 6 (the definition of the kernel size in our proposed method is given in Section 3.3). The results indicate that the configuration $\left\{[17,15,13],[7],[13,7]\right\}$ yields the optimal performance on both image classification and semantic segmentation tasks. Further enlarging the kernels does not lead to additional improvements.
@@
Table G: A comparison of image classification with 384$\times$384 inputs.
@@ Appendix C Speed Analysis
We provide a comparison of speed-accuracy trade-off in Figure 1. More details are listed in Table 1, where an OverLoCK variant often achieves faster speed and higher accuracy simultaneously than a larger variant of another network, demonstrating an excellent trade-off between speed and accuracy. For instance, OverLoCK-XT achieves 1672 imgs/s in throughput, improving upon Swin-T by over 300 imgs/s, while significantly enhancing Top-1 accuracy by 1.4%. Also, OverLoCK-T achieves about 200 imgs/s improvement in throughput compared to ConvNeXt-B while achieving better performance at the cost of only around one-third of the FLOPS. When compared to more advanced models, OverLoCK still exhibits significant advantages. For example, OverLoCK-S surpasses MogaNet-B by over 100 imgs/s in throughput while increasing Top-1 accuracy from 84.3% to 84.8%. Likewise, OverLoCK-XT surpasses BiFormer-T by over 600 imgs/s in throughput while remarkably improving Top-1 accuracy by 1.3%.